OpenStudy (anonymous):

what will an investment of 6000 Euro at 7% p.a compound interest amount to after 5 years if the interest is compounded: a) annually b)quarterly c) monthly? what will an investment of 6000 Euro at 7% p.a compound interest amount to after 5 years if the interest is compounded: a) annually b)quarterly c) monthly? @Mathematics

OpenStudy (anonymous):

A = p(1 + r/n)^(n/t) P = present value R = rate N = compounded how? ( annually = 1, quarterly = 4, monthly = 12, daily = 365 a) A =6000(1 + .07/1)^(1*5)

OpenStudy (anonymous):

Just remember to move your decimal point two to the left when inserting the INTEREST RATE which is R

OpenStudy (anonymous):

A = Final value

OpenStudy (amistre64):

n/t ?

OpenStudy (anonymous):

\[Amount = Principal \times( 1 + \frac{rate}{n})^{nt}\] use this formula for calculating the value when it is annually put n = 1 quaterly then put n= 4 when monthly put n= 12

OpenStudy (anonymous):

my mistake exponent should be multiplied.....thanks for catching that amistre